An electric field with a strength of 1.25x104N/C exerted 2.50 N on a

certain test charge, q. Determine the charge on q.


1
Expert's answer
2022-05-22T14:39:44-0400
"F=qE"

"q=\frac{F}{E}=\frac{2.50}{1.25*10^4}=2.00*10^{-4}\:\rm C"